# Multi-Task Label Discovery via Hierarchical Task Tokens for Partially Annotated Dense Predictions

### v1

@@ Abstract
In recent years, simultaneous learning of multiple dense prediction tasks with partially annotated label data has emerged as an important research area. Previous works primarily focus on constructing cross-task consistency or conducting adversarial training to regularize cross-task predictions, which achieve promising performance improvements, while still suffering from the lack of direct pixel-wise supervision for multi-task dense predictions. To tackle this challenge, we propose a novel approach to optimize a set of learnable hierarchical task tokens, including global and fine-grained ones, to discover consistent pixel-wise supervision signals in both feature and prediction levels. Specifically, the global task tokens are designed for effective cross-task feature interactions in a global context. Then, a group of fine-grained task-specific spatial tokens for each task is learned from the corresponding global task tokens. It is embedded to have dense interactions with each task-specific feature map. The learned global and local fine-grained task tokens are further used to discover pseudo task-specific dense labels at different levels of granularity, and they can be utilized to directly supervise the learning of the multi-task dense prediction framework. Extensive experimental results on challenging NYUD-v2, Cityscapes, and PASCAL Context datasets demonstrate significant improvements over existing state-of-the-art methods for partially annotated multi-task dense prediction.
Keywords:
@@
Multi-task Partially Annotated Dense Prediction Multi-task Learning Label Discovery
1 Introduction
@@
With the rapid development of supervised learning with deep neural networks, various pixel-wise dense prediction tasks with highly complementary properties such as semantic segmentation and depth estimation have achieved great success in multi-task learning (MTL) in recent years misra2016cross ; vandenhende2020mti ; xu2018pad ; ye2022inverted ; zhang2018joint . Researchers pursue the learning of them simultaneously in a unified framework, which can effectively model cross-task correlations and achieve superior results in terms of model training costs and performances.
@@
However, in real-world scenarios, obtaining pixel-level annotations is prohibitively expensive, especially when dealing with a set of distinct dense prediction tasks. Each image has to be annotated with pixel labels for all the tasks. Thus, existing works have delved into the problem of multi-task learning with only partially annotated dense labels li2022learning ; luo2021semi ; wang2022semi ; zamir2020robust ; zeng2019joint . Specifically, as illustrated in Fig. 1 (a), given an input image, for $T$ dense prediction tasks, the task labels are provided for at least one task and at most $T-1$ tasks. Learning a multi-task model under this setting is particularly challenging since every input image lacks some of the task supervision signals, and the performance typically drops significantly if compared to the same model trained with full task label supervisions li2022learning .
@@
Figure 1: (a) Illustration of partially annotated multi-task dense prediction setting. Each input image only has partial task labels from all the tasks. (b) Illustration of the learning of Hierarchical Task Tokens, including global task tokens and fine-grained task tokens, by conducting feature-token interactions in feature and prediction spaces separately. The well-learned hierarchical task tokens can achieve both feature supervision discovery and task label discovery.
@@
Directly discovering pseudo task labels in prediction spaces can alleviate this problem to a certain extent, however, it still suffers from the following two severe limitations: (i) Simply discovering labels in prediction spaces cannot take advantage of the abundant task representations in feature space. For each specific task, labeled and unlabeled images share all of the network parameters during training, and the distributions of their features and predictions should be consistent. For all dense prediction tasks, rich task-generic representations remain in feature space since the backbone parameter sharing, which are beneficial for building cross-task relations and regularizing unsupervised tasks. In this case, excavating labels only in the prediction spaces fails to exploit consistent representations from different levels among different tasks. (ii) Directly cross-task regularizing the prediction space is difficult and expensive. Since representations in the prediction space are highly semantic on each task, additional heavy mapping networks are required to project the different task predictions into a common space to utilize task correlations for learning regularization li2022learning between labeled and unlabeled tasks. Therefore, it is critically important to involve the intermediate task features for the task label discovery and regularization.
To effectively tackle the aforementioned challenges, we propose a novel approach that performs task label discovery from both the feature and prediction spaces via an effective design of learnable Hierarchical Task Tokens (HiTTs). HiTTs are sets of compact parameters that are learned in a hierarchical manner to model global inter-task relationships and local fine-grained intra-task relationships, which allows for discovering pixel-wise task pseudo labels straighforwardly based on the modeled correlations with the task tokens.
@@
More specifically, as depicted in Fig. 1 (b), we apply HiTTs during the multi-task decoding stage and jointly optimize them with the multi-task learning network. The HiTTs consists of two hierarchies. The first hierarchy is a set of global task tokens. The global task tokens are randomly initialized and can perform cross-task feature-token interactions with different task feature maps based on self-attention. These learned task tokens can be used to discover feature-level pseudo supervision by selecting highly activated pixel features correlated to each task. The second hierarchy is the fine-grained task tokens. These tokens are directly derived from the global task tokens with learnable projection layers. They are generated to be spatial token maps to perform interactions within each task-specific feature map at a finer granularity. As the fine-grained task tokens can learn pixel-to-pixel correlation with each task feature map, it thus can help us to discover dense spatial labels for each task. We learn both hierarchies simultaneously in an end-to-end manner, and exploit both levels of supervision signals discovered from the two task-token hierarchies, for optimizing multi-task dense predictions with partially annotated datasets.
@@
  • We propose a novel design of Hierarchical Task Tokens (HiTTs), which can learn hierarchical multi-task representations and correlations with only partially annotated training data.
  • We exploit the hierarchical task tokens to effectively develop multi-task label discovery strategies and obtain direct multi-task supervision signals from both the feature and the prediction levels.
@@
  • Our proposed method significantly outperforms existing state-of-the-art competitors on multi-task partially annotated benchmarks, including NYUD-v2, Cityscapes and PASCAL-Context, and demonstrates clear effectiveness on challenging dense prediction tasks with limited annotations, including segmentation, depth estimation, normal estimation and edge detection, etc.
2 Related Work
Multi-task Dense Prediction. Dense prediction tasks aim to produce pixel-wise predictions for each image. Common tasks including semantic segmentation, depth estimation, and surface normal estimation exhibit high cross-task correlations. For instance, depth discontinuity is usually aligned with semantic boundaries vandenhende2020mti , and surface normal distributions are aligned with spatial derivatives of the depth maps lu2021taskology . Thus, a number of works have been focusing on multi-task dense predictions gao2019nddr ; liu2019end ; misra2016cross ; vandenhende2020mti ; xu2018pad ; yang2023contrastive ; ye2022inverted ; ye2023taskexpert ; taskprompter2023 ; ye2024invpt++ ; zhang2024bridgenet ; zhang2018joint ; zhang2019pattern . They leverage parameters sharing to conduct cross-task interactions by effective attention mechanisms for task feature selection liu2019end , and multi-modal distillation xu2018pad , multi-scale cross-task interactions vandenhende2020mti , global pixel and task interactions ye2022inverted and multi-task mixture-of-experts ye2023taskexpert . However, these works focus on fully-supervised settings. In contrast, our work addresses the challenge of insufficient supervision signals in each task.
Semi-supervised learning. Obtaining pseudo labels for semi-supervised learning is a popular research direction, with several deep learning works published on the topic iscen2019label ; lee2013pseudo ; li2019bidirectional ; shi2018transductive ; sohn2020fixmatch ; tarvainen2017mean ; xie2020self ; zou2018unsupervised ; zou2019confidence . Among them, lee2013pseudo aims at picking up the class which has the maximum predicted confidence. The graph-based label propagation method iscen2019label is also used to infer pseudo-labels for unlabeled data. shi2018transductive provides a confidence level for each unlabeled sample to reduce influences from outliers and uncertain samples, and uses MMF regularization at feature levels to make images with the same label close to each other in the feature space. xie2020self uses accurate pseudo labels produced by the teacher model on clean unlabeled data to train the student model with noise injected. For dense prediction tasks, such as semantic segmentation, several works focus on assigning pixel-wise pseudo annotations from high-confidence predictions li2019bidirectional ; zou2018unsupervised ; zou2019confidence . However, these works target single-task learning setups. Despite pseudo labeling, discovering consistency for regularization is also a popular direction for unlabeled data li2022learning ; luo2021semi ; tang2022towards ; zeng2019joint . li2022learning ; luo2021semi ; zeng2019joint focus on building cross-task consistency, while  tang2022towards uses image-level feature similarities to find important samples for semi-supervised learning. Differently, our work targets pixel-level task supervision discovery by hierarchical task tokens containing multi-level multi-task representations for partially annotated dense predictions.
@@ 2 Related Work
Figure 2: Illustration of our method. HiTTs consist of both global and fine-grained task tokens which learn discriminative task representations by conducting feature-token interactions with attentions in corresponding multi-task decoding stages. The global task tokens $\boldsymbol{\theta_{i}}$ discover feature-level pseudo supervision $\mathcal{L}_{f}$, while the fine-grained task tokens $\boldsymbol{\varphi_{i}}$ inherit the knowledge from global task tokens and directly discover pixel labels for supervision $\mathcal{L}_{p}$. The supervision from ground-truth label is denoted as $\mathcal{L}_{s}$.
@@
Multi-task Partially Supervised Learning. As discussed in the introduction, obtaining pixel-level annotations for every task on images is prohibitively expensive. Therefore, some recent works focus on partially annotated settings for multi-task learning imran2020partly ; li2022learning ; liu2007semi ; lu2021taskology ; luo2021semi ; wang2022semi ; ye2024diffusionmtl ; zamir2020robust ; zeng2019joint . Since directly recovering labels from other tasks is an ill-posed problem li2022learning , enforcing consistency among tasks is usually adopted. For instance, constructing a common feature space to align predictions and impose regularization li2022learning , and leveraging intrinsic connections of different task pairs between predictions of different tasks on unlabeled data in a mediator dataset, when jointly learning multiple models lu2021taskology . Adversarial training is also adopted to align the distributions between labeled and unlabeled data by discriminators wang2022semi . To the best of our knowledge, our hierarchical task tokens for both pseudo feature supervision and task label discovery are a novel exploration of the problem, and show a clear difference from existing works.
3 Proposed Method
@@
Our proposed approach for learning Hierarchical Task Tokens (HiTTs) primarily comprises two stages, i.e., the Global Token Learning and the Fine-grained Token Learning. The overall structure of HiTTs is depicted in Fig. 2. Firstly, in the Global Token Learning stage, the global task tokens produce task features and then learn rich task-level representations by conducting inter- and intra-task interactions with all task features. The global tokens are utilized to exploit rich representations in feature space and discover feature-level pseudo supervision. Subsequently, in the fine-grained stage, we project each task feature into fine-grained feature space by simple convolution layers, and derive the fine-grained tokens from the global tokens by Multi-layer Perceptrons (MLPs), to inherit well-learned global task representations and therefore achieve consistent pseudo label discovery. To perform a uniform confidence-based pseudo label discovery for different types of dense prediction tasks, we follow bruggemann2021exploring to conduct discrete quantization of regression task annotations (e.g. depth estimation and normal estimation), and treat all tasks as pixel-wise classification.
3.1 Global Task Token Learning
@@
In the global task token learning stage, we target learning global tokens representing the distributions of each task, which are further used for pseudo feature supervision discovery. The learning process is mainly achieved by inter- and intra-task interactions among tokens and features, in order to exploit beneficial multi-task representations for token learning and feature supervision discovery.
@@
Given an RGB input $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a multi-task dense prediction framework firstly produces a task-generic representation $\boldsymbol{F_{s}}\in\mathbb{R}^{C\times h\times w}$ through a shared encoder. Considering we have $T$ tasks, and we target decoding task features $\{\boldsymbol{F_{1}},\boldsymbol{F_{2}},\cdots,\boldsymbol{F_{T}}\}$ from $\boldsymbol{F_{s}}$ as well as learning representative global task tokens $\{\boldsymbol{\theta_{1}},\boldsymbol{\theta_{2}},\cdots,\boldsymbol{\theta_{T}}\}$ for each task.

### v2

@@ Abstract.
In recent years, simultaneous learning of multiple dense prediction tasks with partially annotated label data has emerged as an important research area. Previous works primarily focus on leveraging cross-task relations or conducting adversarial training for extra regularization, which achieve promising performance improvements, while still suffering from the lack of direct pixel-wise supervision and extra training of heavy mapping networks. To effectively tackle this challenge, we propose a novel approach to optimize a set of compact learnable hierarchical task tokens, including global and fine-grained ones, to discover consistent pixel-wise supervision signals in both feature and prediction levels. Specifically, the global task tokens are designed for effective cross-task feature interactions in a global context. Then, a group of fine-grained task-specific spatial tokens for each task is learned from the corresponding global task tokens. It is embedded to have dense interactions with each task-specific feature map. The learned global and local fine-grained task tokens are further used to discover pseudo task-specific dense labels at different levels of granularity, and they can be utilized to directly supervise the learning of the multi-task dense prediction framework. Extensive experimental results on challenging NYUD-v2, Cityscapes, and PASCAL Context datasets demonstrate significant improvements over existing state-of-the-art methods for partially annotated multi-task dense prediction.
@@
Multi-task Partially Annotated Dense Prediction, Multi-task Learning, Label Discovery
submissionid: 5591copyright: acmlicensedjournalyear: 2025doi: 10.1145/3746027.3755727conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision tasks
1. Introduction
@@
With the rapid development of supervised learning with deep neural networks, various pixel-wise dense prediction tasks with highly complementary properties such as semantic segmentation and depth estimation have achieved great success in multi-task learning (MTL) in recent years (Misra et al., 2016; Vandenhende et al., 2020; Xu et al., 2018; Ye and Xu, 2022; Zhang et al., 2018). Researchers pursue the learning of them simultaneously in a unified framework, which can effectively model cross-task correlations and achieve superior results in terms of model training costs and performances.
@@
However, in real-world scenarios, obtaining pixel-level annotations is prohibitively expensive, especially when dealing with a set of distinct dense prediction tasks. Each image has to be annotated with pixel labels for all the tasks. Thus, existing works have delved into the problem of multi-task learning with only partially annotated dense labels (Li et al., 2022; Luo et al., 2021; Wang et al., 2022c; Zamir et al., 2020; Zeng et al., 2019; Nishi et al., 2024; Ye and Xu, 2024). Specifically, as illustrated in Fig. 1 (a), given an input image, for $T$ dense prediction tasks, the task labels are provided partially, i.e. for at least one task and at most $T-1$ tasks. Learning a multi-task model under this setting is particularly challenging since every input image lacks some of the task supervision signals, and the performance typically drops significantly if compared to the same model trained with full task label supervisions (Li et al., 2022).
@@
Figure 1. (a) Illustration of partially annotated multi-task dense prediction setting. Each input image only has partial task labels from all the tasks. (b) Illustration of the learning of Hierarchical Task Tokens, including global task tokens and fine-grained task tokens, by conducting feature-token interactions in feature and prediction spaces separately. The well-learned hierarchical task tokens can achieve both feature supervision discovery and task label discovery.
@@
Previous works have been focusing on excavating cross-task relations by training heavy extra mapping networks (Li et al., 2022; Nishi et al., 2024), however, simply applying regularization in compact latent spaces fails to address the lack of dense pixel supervision and limits the performance. On the contrary, directly discovering pseudo task labels in prediction spaces can alleviate this problem to a certain extent, while still suffering from the following two severe limitations: (i) Simply discovering task labels in the prediction spaces separately ignores the highly relevant task relations, thus leading to trivial performance under multitask scenarios. (ii) Solely discovering labels in prediction spaces cannot take advantage of the abundant task representations in feature space. For each specific task, the distributions of the task features and predictions should be consistent. Since the encoder parameters are shared among tasks, thus the produced rich task-generic features are beneficial for building cross-task relations and discovering supervision signals effectively. Therefore, it is critically important to involve hierarchical cross-task representations from feature space to prediction space to consistently boost the task label discovery process.
@@
To effectively tackle the aforementioned challenges, we propose a novel approach that performs task label discovery from both the feature and prediction spaces via an effective design of learnable Hierarchical Task Tokens (HiTTs). HiTTs are sets of compact parameters that are learned in a hierarchical manner to model global inter-task relationships and local fine-grained intra-task relationships, which allows for discovering pixel-wise task pseudo labels straightforwardly in both feature and prediction spaces consistently. More specifically, as depicted in Fig. 1 (b), we apply HiTTs during the multi-task decoding stage and jointly optimize them with the multi-task learning network. The HiTTs consist of two hierarchies. The first hierarchy is a set of global task tokens. The global task tokens are randomly initialized and can perform cross-task feature-token interactions with different task feature maps based on self-attention. These learned task tokens can be used to discover feature-level pseudo supervision by selecting highly activated pixel features correlated to each task. The second hierarchy is the fine-grained task tokens. These tokens are directly derived from the global task tokens with learnable projection layers to inherent beneficial task representations. They are subsequently utilized to perform interactions within each task-specific feature map at a finer granularity. As the fine-grained task tokens can learn pixel-to-pixel correlation with each task feature map, they thus benefit the discovery of dense spatial labels for each task. Compared with naive pseudo labeling process (Lee et al., 2013), the HiTTs can bridge effective information from supervised tasks so as to encourage highly confident predictions on unsupervised tasks. We learn both hierarchies of tokens simultaneously in an end-to-end manner incorporating the multi-task baseline network, and exploit both levels of supervision signals discovered from the two hierarchies, for optimizing multi-task dense predictions on partially annotated datasets.
@@
  • Instead of discovering cross-task regularization by extra heavy mapping networks, we propose to utilize cross-task relations for high-quality task label discovery, which serves as pixel-level dense pseudo supervision under the multi-task partially supervised setting.
  • We propose a novel design of Hierarchical Task Tokens (HiTTs), which can learn hierarchical multi-task representations for high-quality pseudo label discovery consistently in both the feature and the prediction levels.
@@
  • Our proposed method significantly outperforms existing state-of-the-art competitors on multi-task partially annotated benchmarks, including PASCAL-Context, NYUD-v2 and Cityscapes, and demonstrates clear effectiveness on challenging dense prediction tasks with limited annotations, including segmentation, depth estimation, normal estimation and edge detection, etc. Code is released at https://github.com/Evergreen0929/EEMTL.
@@
Figure 2. Illustration of our method. HiTTs consist of both global and fine-grained task tokens which learn discriminative task representations by conducting feature-token interactions with attentions in corresponding multi-task decoding stages. The global task tokens $\boldsymbol{\theta_{i}}$ discover feature-level pseudo supervision $\mathcal{L}_{f}$, while the fine-grained task tokens $\boldsymbol{\varphi_{i}}$ inherit the knowledge from global task tokens and directly discover pixel labels for supervision $\mathcal{L}_{p}$. The supervision from ground-truth label is denoted as $\mathcal{L}_{s}$.
2. Related Work
Multi-task Dense Prediction. Dense prediction tasks aim to produce pixel-wise predictions for each image. Common tasks including semantic segmentation, depth estimation, and surface normal estimation exhibit high cross-task correlations. For instance, depth discontinuity is usually aligned with semantic boundaries (Vandenhende et al., 2020), and surface normal distributions are aligned with spatial derivatives of the depth maps (Lu et al., 2021). Thus, a number of works have been focusing on multi-task dense predictions (Gao et al., 2019; Liu et al., 2019; Misra et al., 2016; Vandenhende et al., 2020; Xu et al., 2018; Yang et al., 2023b; Ye and Xu, 2022, 2023a, 2023b; Zhang et al., 2023, 2018, 2019, 2025; Tang et al., 2025; Tian et al., 2024; Lu et al., 2024; Cao et al., [n. d.]; Yang et al., [n. d.]; Chavhan et al., 2025). They leverage parameters sharing to conduct cross-task interactions by effective attention mechanisms for task feature selection (Liu et al., 2019), and multi-modal distillation (Xu et al., 2018), multi-scale cross-task interactions (Vandenhende et al., 2020), global pixel and task interactions (Ye and Xu, 2022) and multi-task bridge features (Zhang et al., 2025). However, these works focus on fully-supervised settings. In contrast, our work addresses the challenge of insufficient supervision signals in each task.
Semi-supervised learning. Obtaining pseudo labels for semi-supervised learning is a popular research direction, with several deep learning works published on the topic (Iscen et al., 2019; Lee et al., 2013; Li et al., 2019; Shi et al., 2018; Sohn et al., 2020; Tarvainen and Valpola, 2017; Xie et al., 2020; Zou et al., 2018, 2019). Among them, (Lee et al., 2013) aims at picking up the class which has the maximum predicted confidence. The graph-based label propagation method (Iscen et al., 2019) is also used to infer pseudo-labels for unlabeled data. (Shi et al., 2018) provides a confidence level for each unlabeled sample to reduce influences from outliers and uncertain samples, and uses MMF regularization at feature levels to make images with the same label close to each other in the feature space. (Xie et al., 2020) uses accurate pseudo labels produced by the teacher model on clean unlabeled data to train the student model with noise injected. For semi-supervised dense prediction tasks (Pieropan et al., 2022; Yang et al., 2023a; Hoyer et al., 2024; Mai et al., 2024; Ran et al., 2024; Yang et al., 2025), such as semantic segmentation, several works focus on assigning pixel-wise pseudo annotations from high-confidence predictions (Li et al., 2019; Zou et al., 2018, 2019). However, these works target single-task learning setups. Despite pseudo labeling, discovering consistency for regularization is also a popular direction for unlabeled data (Li et al., 2022; Luo et al., 2021; Tang and Jia, 2022; Zeng et al., 2019). (Li et al., 2022; Luo et al., 2021; Zeng et al., 2019) focus on building cross-task consistency, while  (Tang and Jia, 2022) uses image-level feature similarities to find important samples for semi-supervised learning. Differently, our work targets pixel-level task supervision discovery by hierarchical task tokens containing multi-level multi-task representations for partially annotated dense predictions.
@@ 2. Related Work
Multi-task Partially Supervised Learning. As discussed in the introduction, obtaining pixel-level annotations for every task on images is prohibitively expensive. Therefore, some recent works focus on partially annotated settings for multi-task learning (Imran et al., 2020; Li et al., 2022; Liu et al., 2007; Lu et al., 2021; Luo et al., 2021; Wang et al., 2022c; Ye and Xu, 2024; Zamir et al., 2020; Zeng et al., 2019; Nishi et al., 2024). Since directly recovering labels from other tasks is an ill-posed problem (Li et al., 2022), enforcing consistency among tasks is usually adopted. For instance, constructing a common feature space to align predictions and impose regularization (Li et al., 2022), and leveraging intrinsic connections of different task pairs between predictions of different tasks on unlabeled data in a mediator dataset, when jointly learning multiple models (Lu et al., 2021). Adversarial training is also adopted to align the distributions between labeled and unlabeled data by discriminators (Wang et al., 2022c), and multi-task denoising diffusion is adopted (Ye and Xu, 2024) to address the issue of noise in initial prediction maps. To the best of our knowledge, our hierarchical task tokens for both pseudo feature supervision and task label discovery are a novel exploration of the problem, and show a clear difference from existing works.
3. Proposed Method
@@
Our proposed approach for learning Hierarchical Task Tokens (HiTTs) primarily comprises two stages, i.e., the Global Token Learning and the Fine-grained Token Learning. The overall structure of HiTTs is depicted in Fig. 2. Firstly, in the Global Token Learning stage, the global task tokens produce task features and then learn rich task-level representations by conducting inter- and intra-task attention with all task features. The global tokens are utilized to exploit rich representations in feature space and discover feature-level pseudo supervision. Subsequently, in the fine-grained stage, we project each task feature into fine-grained feature space by simple convolution layers, and derive the fine-grained tokens from the global tokens by Multi-layer Perceptrons (MLPs), to inherit well-learned global task representations and therefore achieve consistent pseudo label discovery. To perform a uniform confidence-based pseudo label discovery for different types of dense prediction tasks, we follow (Brüggemann et al., 2021) to conduct discrete quantization of regression task annotations (e.g. depth estimation and normal estimation), and treat all tasks as pixel-wise classification.
Figure 3. (a) Global Task Token Learning. It mainly contains two stages: Inter- and Intra-Task Attention for thoroughly feature-token cross-task interaction to obtain robust and representative global task tokens, and excavating pseudo feature supervision based on the learned task tokens. (b) Fine-grained Task Token Learning. The projected fine-grained tokens and feature maps (projection is shown in Fig. 4) can be used for discovering high-quality pseudo labels in prediction spaces.
3.1. Global Task Token Learning
@@
In the global task token learning stage, we target learning global tokens representing the distributions of each task, which are further used for pseudo feature supervision discovery. The learning process is mainly achieved by inter- and intra-task attention among tokens and features, in order to exploit beneficial multi-task representations for token learning and feature supervision discovery.
@@
Given an RGB input $\boldsymbol{X}\in\mathbb{R}^{3\times H\times W}$, a multi-task dense prediction framework firstly produces a task-generic representation $\boldsymbol{F_{s}}\in\mathbb{R}^{C\times h\times w}$ through a shared encoder. Considering we have $T$ tasks, and we target decoding task features $\{\boldsymbol{F_{1}},\boldsymbol{F_{2}},\cdots,\boldsymbol{F_{T}}\}$ from $\boldsymbol{F_{s}}$ as well as learning representative global task tokens $\{\boldsymbol{\theta_{1}},\boldsymbol{\theta_{2}},\cdots,\boldsymbol{\theta_{T}}\},\boldsymbol{\theta_{i}}\in\mathbb{R}^{C}$ for each task. They are randomly initialized learnable vectors serving as additional input tokens for the decoder.